# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USAthanks: Present address: Microsoft Azure Quantum, Redmond, WA, USA

### v1

@@ II Results
Finally, we measure the fraction of bit-flip errors while performing the CZ gate. From [33], we expect idling bit-flip errors to be small and here we measure the added bit-flip error from executing the CZ gate. We perform the following measurement protocol for measuring bit-flip errors on the control (target) qubit. We initialize the control (target) in a basis state, |$\ket{01}$ or |$\ket{10}$, and perform a Ramsey-style circuit on the target (control) qubit with $N$ CZ gates sandwiched between the $\pi/2$ pulses. We measure the fraction of bit-flip outcomes on the control (target) qubit. These results are summarized in Fig. 3d-f. As apparent from the figure, the overall probability of bit flips is bounded at few parts per million per gate, probably limited by leakage misassignments. We note that although this particular gate preserves the Pauli noise-bias, we do not anticipate using a full gate set that preserves Pauli noise-bias.
Figure 4: Leakage propagation properties of the CZ gate. a. Leakage on the control qubit to |$\ket{00}$ prior to the CZ gate causes the gate to be switched off and instead the identity process is performed on both qubits. b. If photon loss occurs during the gate, the error channel can be modeled as either leakage occurring before (and the gate is switched off as above) or after the CZ gate (resulting in a CZ error). c. Circuit diagram used to test the leakage propagation properties of the CZ gate. For different preparations on the control qubit, quantum process tomography is performed on the target qubit after applying $N$ CZ gates on the target qubit. To test for control qubit leakage properties, the tomography is conditioned on the control qubit being found in the leakage |$\ket{00}$ state. d.-f. Conditioned process tomography on target qubit when the control qubit is prepared in |$\ket{00}$, |$\ket{10}$, and |$\ket{01}$, respectively. For each case, the real (upper) and imaginary (lower) part of the $\chi$-error matrices are plotted for $N=1$ CZ gates (left) and $N=27$ CZ gates (right). For control qubit initialized in |$\ket{00}$ and |$\ket{10}$, the expected identity process is observed indicating that a leakage error turns off the gate; while the expected ‘CZ error’ is observed for leakage when the control qubit is initialized in |$\ket{01}$ indicated by large diagonal components for $I$ and $Z$ in the real part of $\chi$-error matrix.
@@
In our system, we consider two possible leakage scenarios: photon loss to |$\ket{00}$ that occurs prior to a CZ gate and leakage that occurs mid-way through the CZ gate at unknown time. Furthermore, we study the more likely scenario of leakage affecting only the control qubit, although due to the symmetry of a CZ gate our results are expected to apply similarly to the case of leakage on the target qubit. Whilst other dual-rail leakage states are possible due to cavity heating, such as |$\ket{02}$ or |$\ket{11}$, these occur at least $1000$ times less frequently in our hardware and are not considered here.
To see why leakage propagation is particularly benign, we can first consider what happens if the control qubit is in the |$\ket{00}$ state prior to the CZ gate. Since there is only one excitation in the system (in the target qubit), the dispersive interaction is never activated resulting in an identity operation on the target qubit. The exact same scenario occurs if the control qubit begins in |$\ket{10}$ and suffers from photon loss during the CZ gate: since no photons were ever swapped into the coupler, the dispersive interaction was never activated which again results in an identity operation on the target qubit. The interesting case occurs when the control qubit begins in |$\ket{01}$ and swaps into the coupler. Should excitation loss occur at any time during the gate, the result will be an incomplete CPHASE gate with unknown rotation angle. Averaging over all possible loss times results in what we term a ‘CZ error’, a correlated error in which the target qubit dephases, but only if the control leaks from the initial |$\ket{01}$ state. Such an error model is useful for enabling delayed erasure checks. A more detailed and theoretical treatment of the leakage error model can be found in Appendix E.
@@
We first consider the two leakage cases when initializing the control qubit in the |$\ket{00}$ and |$\ket{10}$ state. Our results, shown in Fig. 4d,e for $N=1$ and $N=27$ CZ gates, show the expected behavior, that the resulting process on the target qubit is close to the identity process. We use the $\chi$-error matrix and quantify the deviation from the identity process by extracting the residual total Pauli error (add definition). For $N=1$, we extract a small residual Pauli error of $0.19(2)\%$ and $0.18(2)\%$ for |$\ket{00}$ and |$\ket{10}$ preparation, respectively. By repeating the gate many times, we are able to increase the fraction of erasures that occurred during the CZ gates over those that occurred during SPAM. At $N=27$ we estimate the fraction of CZ erasures to be roughly $2/3$ of total erasures. We extract a still small residual Pauli error of $0.33(2)\%$ and $0.35(2)\%$, experimentally confirming that these types of leakage errors essentially turn off the gate.
Finally, we consider the case when the control qubit is in |$\ket{01}$ and the CZ interaction is turned on for a random duration until the leakage event turns off the interaction. The resulting process on the target qubit is shown in Fig. 4f, and closely matches a pure dephasing channel. For $N=1$ gates the extracted residual $X$ and $Y$ errors total $0.7(1)\%$ and at $N=27$ this error totals $2.55(5)\%$, a number that quantifies the deviation of this error channel from the theoretically expected dephasing channel. At $N=27$ assuming all of the error comes from the CZ gates, we find an error of $0.094(2)\%$ per gate, which quantifies the deviation of our target qubit error channel from our ‘CZ error’ channel. When combined with the small fraction of leakage per CZ gate of $0.53\%$, this error (which can be thought of as an additional contribution to our Pauli error in a QEC setting) drops to the order $10^{-5}$ or lower.
@@ III Discussion and conclusion
In summary, we have devised and demonstrated what is to our knowledge the first entangling gate for superconducting erasure qubits. This SWS gate uses the existing transmon parametric couplers to allow sub-microsecond gate times, and achieves state-of-the-art performance. Moreover, it preserves (see Table 1) the desirable error hierarchy and long coherence times of the dual-rail cavity qubit. By virtue of being an excitation preserving operation, the dominant errors ($\sim 0.5\%$) remain erasures which can be detected by end-of-the-line [33] or mid-circuit [36, 37] erasure check measurements. The high post-selected fidelity (99.97-99.90%) is then primarily limited by dephasing in the dual-rail and the transmon coupler, the strong bias against bit flips is maintained, which are bounded by our measurements to be less than a few parts per million per gate. The gate displays a significant asymmetry in the errors induced on the control (i.e. the dual-rail that swaps into the coupler) and the target qubits. Remarkably, we observe that the Pauli error-per-gate on the target qubit is as low as 0.01%. Finally, the gate has a very simple behavior when one of the qubits is already erased or suffers an erasure during the gate, which should minimize the propagation of errors and reduces the frequency of mid-circuit erasure checks that will be needed. Further improvements in performance by increasing the gate speed or the coherences of the coupler and cavity dual-rails may be possible in future implementations.
@@
The addition of a fast and high-fidelity entangling operation completes the gate set for universal operations for the dual-rail cavity qubit, and can enable high performance in short-depth circuits and new types of error mitigation. In addition, the properties of SWS gate are highly optimal in the context of quantum error correction. For example, the erasure and dephasing rates are a factor of ten below their respective separate thresholds for the surface code. In the presence of both kinds of errors, detailed numerical simulations [40] are needed to estimate the actual performance but the results described here provide a realistic way to realize large error correction gains of 10 or more with increase in code distance in the near future, making the superconducting dual-rail cavity qubit an attractive path for achieving fault-tolerant quantum computing.
Acknowledgments
We would like to thank the mechanical engineering, control systems, fabrication, and software teams, as well as the rest of the team at Quantum Circuits, Inc. for their contributions to the success of this project. In particular, we would like to recognize crucial efforts from Beau Birdsall, Hannah Hastings, and Jonathan Strenczewilk for contributions towards mechanical design, hardware assembly, and deployment; Michael Maxwell for leading control system deployments for this project; Amos Anderson, Luke Mastalli-Kelly, Jonathan Chen, Doug Szperka for developing software used for this project. We would also like to thank Luigi Frunzio, Sophia Xu, John Garmon, and Takahiro Tsunoda for their feedback on the manuscript.
Author contributions
@@ Acknowledgments
J.D.T. and S.O.M. contributed to the proposal and concept of the SWS gate. S.O.M. led the design of the dual-rail hardware architecture. J.O.Y led the hardware deployment and experimental setup for this project. C.U.L. led the cavity design, preparation, and testing. K.S.C. led two-qubit gate testing, measurements, and analysis. N.M., T.N., K.S.C, P.L., B.H.E., A.A., A.N. and J.D.T. contributed to data acquisition and data analysis in all experiments. T.N., N.M., P.L., A.A., and J.D.T. contributed to the SWS gate calibration procedure. J.D.T. and K.S.C. contributed to numerical simulations of the SWS gate. P.L., N.M., J.O.Y., T.N., A.A., K.S.C., G.L., and N.K. contributed to system bring-up and system calibration. S.O.M., J.O.Y, R.D.L., T.Z.C., H.M., R.C., S.M.F. and D.K.W. contributed to quantum hardware design and simulation. S.O.M. and R.C. contributed to coupler design and flux-pump delivery. C.U.L., C.K., J.O.Y., B.H.E. and N.K. contributed to cavity preparation, device testing and assembly. S.S and B.G. contributed to device fabrication. J.O.Y., S.O.M., G. L., N.K. and R.A.S. contributed to RF wiring and implementation of the readout chain. J.C.C., P.L., N.M., T.K. contributed to data analysis of randomized benchmarking results and supporting simulations. J.D.T., S.O.M., A.N, N.M., K.S.C. and T.S. contributed to an earlier version of the SWS gate. K.S.C, R.J.S., N.M., T.N. and J.D.T. wrote the manuscript. R.J.S., S.H.M., J.A. supervised the project. All authors contributed to editing and reviewing the manuscript.
Competing interests statement
@@
Appendix A Two-qubit gate calibration
In this section, we discuss the calibration of the CZ gate in detail. The calibration procedure is illustrated in the flowchart in Fig.  5a, which starts with calibrating the photon swap operation between the control qubit cavity $a_{2}$ and the coupler $c$, referred to as the cavity-coupler swap. The underlying interaction for this cavity-coupler swap is identical to that used for the single-qubit gate. However, in this case, the parametric flux pump is applied at half the frequency difference of the Stark shifted SQUID-transmon coupler and cavity $a_{2}$. A coherent transfer of a single photon between the cavity and the coupler is shown in Fig. 5b where we also present the stroboscopic evolution of the photon population in the cavity $P_{\rm{cav}}$ at the resonant parametric pump frequency, which follows
$P_{\rm{cav}}=\frac{1}{2}e^{-{\kappa t}}\left(1+e^{-\kappa_{\phi}t}{\rm{cos}}(g_{ac}t)\right).$ (2)
Here, $g_{ac}$ represents the strength of the cavity-coupler swap, while $\kappa$ and $\kappa_{\phi}$ denote the photon decay rate and the dephasing rate when the parametric drive is turned on. As the first step in calibrating the cavity coupler swap operation, we find the amplitude $A_{p}$ that maximizes the metric $g_{ac}/\kappa_{\phi}$, which leads to the highest fidelity of the cavity-coupler swap operation when you postselect for photon loss. Following this, we calibrate the duration of the swap pulse $t_{\mathrm{SWAP}}$ by repeatedly applying the pulse after starting with a photon in cavity $a_{2}$ and minimizing the population in the cavity $a_{2}$ at the end of odd number of swap operations (see Fig. 5c). When the swap operation is used in the SWS gate sequence, an additional parameter needs to be specified which is the relative phase difference between the first and the second swap operations. To understand why this is important, we first note that at the end of the CZ gate we want all the population in cavity $a_{2}$ to return to its original value irrespective of which two qubit state we start in. This condition is trivially satisfied for starting states |$\ket{10g01},\ket{10g10}$ and |$\ket{01g01}$ since in these cases either no transfer of population takes place between cavity $a_{2}$ and coupler $c$ or an on-resonant swap takes place, either way the phase of the second swap operation is irrelevant. The only non-trivial case occurs when we start in the state |$\ket{01g10}$, in which case we perform a swap operation detuned by $\chi_{bc}$ (see Fig. 11). In this case we can close the Bloch sphere trajectory by choosing the correct phase of the second swap operation in the gate. To calibrate this parameter we initialize our system in the state |$\ket{01g10}$ and apply the SWS gate sequence while sweeping the phase of the second swap pulse and measure the total erasure fraction of the circuit. We choose the swap back phase $\phi_{\mathrm{swap}}$ that minimizes the erasure fraction. Since our gate is excitation-preserving, every time we leave population in the coupler mode $c$ we will find the control qubit in the vacuum state. Therefore, minimizing the erasure fraction is equivalent to maximizing the population in cavity $a_{2}$ at the end of the gate. It is important to note that as the delay time varies, the swap back phase should be calibrated and adjusted accordingly prior to measuring the corresponding entangling phase $\phi_{e}$ as illustrated in the flowchart, until the delay time is accurately calibrated.
Figure 5: Calibration of CZ gate. a Flowchart of the CZ gate calibration procedure. b Left: Schematic illustration of swap between the control qubit cavity $a_{2}$ (red) and coupler $c$ (purple). Right: Photon population in the control qubit cavity as a function of parametric pump frequency $\omega_{p}$ and duration of the pump, and additionally the stroboscopic evolution at the resonant pump frequency marked by a red dashed line in a. c-f. Each Fig. shows the pulse sequence and the corresponding calibration result. c. Calibration of duration of cavity-coupler swap by measuring the population in the control qubit cavity after odd numbers of swap. d. Calibration of the phase of rotation for the second swap operation in the CZ gate by minimizing the erasure fraction. e. Calibration of the delay time of the CZ gate by measuring the entangling phase per CZ gate, which should corresponds to $\pi$ for the calibrated delay time marked by dashed line. f. Calibration of single-qubit Z rotations per CZ gate. The Ramsey-like protocol should be performed on either the control or target qubit to measure Z rotation of each qubit.
With the cavity-coupler swap operation calibrated, both the duration of the pulse and the swap back phase for a given delay time $t_{\mathrm{wait}}$, the unitary of the SWS gate can be written as the following,
where $\phi_{e}$ is the entangling phase, $\phi_{a}$ and $\phi_{b}$ are respective $Z$-phase on each qubit accumulated as a result of the gate. In the ideal case of instantaneous cavity-coupler swap, the entangling phase is given by $\phi_{e}=\chi_{bc}\cdot t_{\mathrm{wait}}$ where $\chi_{bc}$ is the cross-Kerr between the coupler and target qubit cavity $b_{1}$. However, in practice, the parameters associated with the implementation of gate such as a finite duration and ramp time of cavity-coupler swap as well as the presence of cross-Kerr $\chi_{ab}$ between control and target qubit cavities should be taken into account. The origin of the Z-phase on the qubits includes the Stark shift during the swap operations which affects both control and target qubits, and photon population leaving the qubit manifold which only affects the control qubit whose constituent cavity takes part in the population transfer into the coupler during the gate. In order to realize a CZ gate, it is critical to calibrate the delay time such that entangling phase $\phi_{e}$ corresponds to $\pi$, followed by the calibration of $\phi_{a}$ and $\phi_{b}$.
Calibration of the delay time can be done by using two SWS gates separated by $X$ gate on both qubits [45], which can be described by the following unitary
$\begin{split}&U_{\rm{echoed}}\\ &=\begin{pmatrix}0&0&0&e^{i(\phi_{a}+\phi_{b}+\phi_{e})}\\ 0&0&e^{i(\phi_{a}+\phi_{b})}&0\\ 0&e^{i(\phi_{a}+\phi_{b})}&&0\\ e^{i(\phi_{a}+\phi_{b}+\phi_{e})}&0&0&0\end{pmatrix}.\end{split}$ (4)
This echoed gate sequence has an advantage that it does not require the calibration of $\phi_{a}$ and $\phi_{b}$ to extract the entangling phase $\phi_{e}$ for a given delay time. To be more specific, as depicted in Fig. 5e, if we prepare a superposition state by applying $X_{\pi/2}$ gate on the control qubit, i.e., $\frac{1}{\sqrt{2}}(|0_{L}0_{L}\rangle-i|1_{L}0_{L}\rangle)$ where the logical basis states are defined by |$\ket{0_{L}}=\ket{01}$ and |$\ket{1_{L}}=\ket{10}$ as in the main text, the above unitary $U_{\rm{echoed}}$ results in a relative phase shift $\phi_{e}$ in the superposition state, $\frac{1}{\sqrt{2}}(e^{i\phi_{e}}|1_{L}1_{L}\rangle-i|0_{L}1_{L}\rangle)$, with $\phi_{a}+\phi_{b}$ factored out as a global phase shift. In order to amplify the sensitivity of the entangling phase to the delay time, we can apply even number of multiple SWS gates, $N$ in total, where the first $N/2$ gates and the second $N/2$ gates are separated by $X_{\pi}$ gate on both qubits, which leads to $N\phi_{e}/2$ of phase shift in the superposition state. We can repeatedly perform this protocol until we find the optimal delay time that leads to $\phi_{e}=\pi$. It should be reminded that for each delay time the swap back phase in the SWS gate should be calibrated prior to running the protocol to measure $\phi_{e}$. The entangling phase as a function of delay time is shown for different numbers of SWS gates applied in the calibration sequence.
Once the delay time is precisely calibrated, we now calibrate the Z-phase on each qubit by using similar protocols. As shown in Fig. 5f, application of SWS gate leads to a phase shift of either $\phi_{a}$ or $\phi_{b}$ in the superposition state prepared by applying $X_{\pi/2}$ gate on either control or target qubit. Similarly to the case of entangling phase, we can amplify the phase shift by applying multiple SWS gates, $N$ in total, which leads to $N\phi_{a}$ or $N\phi_{b}$ of phase shift in the superposition state.
Appendix B CZ gate on-off ratio

### v2

@@ II Results
Finally, we measure the fraction of bit-flip errors while performing the CZ gate. From [34], we expect idling bit-flip errors to be small and here we measure the added bit-flip error from executing the CZ gate. We perform the following measurement protocol for measuring bit-flip errors on the control (target) qubit. We initialize the control (target) in a basis state, |$\ket{01}$ or |$\ket{10}$, and perform a Ramsey-style circuit on the target (control) qubit with $N$ CZ gates sandwiched between the $\pi/2$ pulses. We measure the fraction of bit-flip outcomes on the control (target) qubit. These results are summarized in Fig. 3d-f. As apparent from the figure, the overall probability of bit flips is bounded at few parts per million per gate, probably limited by leakage misassignments. We note that although the CZ gate preserves the Pauli noise-bias, we do not anticipate using a full gate set that preserves Pauli noise-bias.
Figure 4: Leakage propagation properties of the CZ gate. a. Leakage on the control qubit to |$\ket{00}$ prior to the CZ gate causes the gate to be switched off and instead the identity process is performed on both qubits. b. If photon loss occurs during the gate, the error channel can be modeled as either leakage occurring before (and the gate is switched off as above) or after the CZ gate (resulting in a CZ error). c. Circuit diagram used to test the leakage propagation properties of the CZ gate. For different preparations on the control qubit, quantum process tomography is performed on the target qubit after applying $N$ CZ gates on the target qubit. To test for control qubit leakage properties, the tomography is conditioned on the control qubit being found in the leakage |$\ket{00}$ state. d.-f. Conditioned process tomography on target qubit when the control qubit is prepared in |$\ket{00}$, |$\ket{10}$, and |$\ket{01}$, respectively. For each case, the real (upper) and imaginary (lower) part of the $\chi$-error matrices are plotted for $N=1$ CZ gates (left) and $N=27$ CZ gates (right). The color scale is clipped at 0.1 to highlight deviations from the expected error channel. For control qubit initialized in |$\ket{00}$ and |$\ket{10}$, the expected identity process is observed indicating that a leakage error turns off the gate; while the expected ‘CZ error’ is observed for leakage when the control qubit is initialized in |$\ket{01}$ indicated by large diagonal components for $I$ and $Z$ in the real part of $\chi$-error matrix.
@@
In our system, we consider two possible leakage scenarios: photon loss to |$\ket{00}$ that occurs prior to a CZ gate and leakage that occurs mid-way through the CZ gate at an unknown time. Furthermore, we study the more likely scenario of leakage affecting only the control qubit, although due to the symmetry of a CZ gate our results are expected to apply similarly to the case of leakage on the target qubit. Whilst other dual-rail leakage states are possible due to cavity heating, such as |$\ket{02}$ or |$\ket{11}$, these occur at least $1000$ times less frequently in our hardware and are not considered here.
To see why leakage propagation is particularly benign, we can first consider what happens if the control qubit is in the |$\ket{00}$ state prior to the CZ gate. Since there is only one excitation in the system (in the target qubit), the dispersive interaction is never activated resulting in an identity operation on the target qubit. The exact same scenario occurs if the control qubit begins in |$\ket{10}$ and suffers from photon loss during the CZ gate: since no photons were ever swapped into the coupler, the dispersive interaction was never activated which again results in an identity operation on the target qubit. The interesting case occurs when the control qubit begins in |$\ket{01}$ and swaps into the coupler. Should excitation loss occur at any time during the gate, the result will be an incomplete CPHASE gate with unknown rotation angle. Averaging over all possible loss times results in what we term a ‘CZ error’, a correlated error in which the target qubit dephases, but only if the control leaks from the initial |$\ket{01}$ state. Such an error model is useful for enabling delayed erasure checks. A more detailed and theoretical treatment of the leakage error model can be found in Appendix F.
@@
We first consider the two leakage cases when initializing the control qubit in the |$\ket{00}$ and |$\ket{10}$ state. Our results, shown in Fig. 4d,e for $N=1$ and $N=27$ CZ gates, show the expected behavior, that the resulting process on the target qubit is close to the identity process. We use the $\chi$-error matrix and quantify the deviation from the identity process by extracting the residual total Pauli error. For $N=1$, we extract a small residual Pauli error of $0.19(2)\%$ and $0.18(2)\%$ for |$\ket{00}$ and |$\ket{10}$ preparation, respectively. By repeating the gate many times, we are able to increase the fraction of erasures that occurred during the CZ gates over those that occurred during SPAM. At $N=27$ we estimate the fraction of CZ erasures to be roughly $2/3$ of total erasures. We extract a still small residual Pauli error of $0.33(2)\%$ and $0.35(2)\%$, experimentally confirming that these types of leakage errors essentially turn off the gate.
Finally, we consider the case when the control qubit is in |$\ket{01}$ and the CZ interaction is turned on for a random duration until the leakage event turns off the interaction. The resulting process on the target qubit is shown in Fig. 4f, and closely matches a pure dephasing channel. For $N=1$ gates the extracted residual $X$ and $Y$ errors total $0.7(1)\%$ and at $N=27$ this error totals $2.55(5)\%$, a number that quantifies the deviation of this error channel from the theoretically expected dephasing channel. Dividing this number by $N$ yields $0.094(2)\%$, a number that characterizes an upper bound for the $X$ and $Y$ Pauli errors experienced by the target qubit when the control qubit leaks during a CZ gate.
@@ III Discussion and conclusion
In summary, we have devised and demonstrated what is to our knowledge the first entangling gate for superconducting erasure qubits. This SWS gate uses the existing transmon parametric couplers to allow sub-microsecond gate times, and achieves state-of-the-art performance. Moreover, it preserves (see Table 1) the desirable error hierarchy and long coherence times of the dual-rail cavity qubit. By virtue of being an excitation preserving operation, the dominant errors ($\sim 0.5\%$) remain erasures which can be detected by end-of-the-line [34] or mid-circuit [37, 38] erasure check measurements. The high postselected fidelity (99.97-99.90%) is then primarily limited by dephasing in the dual-rail and the transmon coupler, the strong bias against bit flips is maintained, which are bounded by our measurements to be less than a few parts per million per gate. The gate displays a significant asymmetry in the errors induced on the control (i.e. the dual-rail that swaps into the coupler) and the target qubits. Remarkably, we observe that the Pauli error-per-gate on the target qubit (after postselection) is as low as 0.01%. Finally, the gate has a very simple behavior when one of the qubits is already erased or suffers an erasure during the gate, which should minimize the propagation of errors and reduces the frequency of mid-circuit erasure checks. Further improvements in performance by increasing the gate speed or the coherences of the coupler and dual-rails may be possible in future implementations.
@@
The addition of a fast and high-fidelity entangling operation completes the gate set for universal operations for the dual-rail cavity qubit, and can enable high performance in short-depth circuits and new types of error mitigation. In addition, the properties of SWS gate are highly optimal in the context of quantum error correction. For example, the erasure and dephasing rates are a factor of ten below their respective separate thresholds for the surface code. In the presence of both kinds of errors, detailed numerical simulations [41] are needed to estimate the actual performance, but the results described here provide a realistic way to realize large error correction gains of 10 or more with increase in code distance in the near future, making the superconducting dual-rail cavity qubit an attractive path for achieving fault-tolerant quantum computing.
IV Acknowledgments
We would like to thank Luigi Frunzio, Sophia Xue, John Garmon, and Takahiro Tsunoda for their feedback on the manuscript. We would also like to acknowledge Ray Smets for his executive leadership and support for the entire Quantum Circuits team.
V Author list
Quantum Circuits, Inc. Team and Collaborators
Nitish Mehta1,*, James D. Teoh1,*, Taewan Noh1,*, Ankur Agrawal1, Amos Anderson1, Beau Birdsall1, Avadh Brahmbhatt1, Winfred Byrd1, Anthony Cabrera1, Marc Cacioppo1, Leo Carroll1, Jonathan Chen1, Tzu-Chiao Chien1,4, Richard Chamberlain1, Jacob C. Curtis1, Doreen Danso1, Sanjana Renganatha Desigan1, Francesco D’Acounto1, Bassel Heiba Elfeky1, S. M. Farzaneh1, Chase Foley1, Benjamin Gudlewski1, Hannah Hastings1, Robert Johnson1, Nishaad Khedkar1,2,3, Trevor Keen1, Anup Kumar1, Cihan Kurter1, Kamila Krawczuk1, Eric Langstengel1, Richard D. Li1, Gangqiang Liu1, Hanyi Lu1, Pinlei Lu1, Luke Mastalli-Kelly1, Adam Maines1, Michael Maxwell1, Heather McCarrick1, Mona Mirzaei1, Anirudh Narla1,4, Omar Rashad1, Erik Reikes1, Mizanur Rahman1, Rurik Primiani1, Michael Schwaller1, Ali Sabbah1, Tali Shemma1,5, Ruby A. Shi1, Sitakanta Satapathy1, Dean Stolpe1, Jonathan Strenczewilk1, Doug Szperka1, Iu-Wei Sze1, David Sweeney1, Preetham Tikkireddi1, Chin-Lun Tsung1, Daren Vet Sam1, Daniel K. Weiss1, Zhibo Yang1, Liuqi Yu1, Teng Zhang1, Olivier Boireau1, Stephen Horton1, Sean Weinberg1, José Aumentado1, Bryan Cord1, Chan U Lei1, Joseph O. Yuan1, Shantanu O. Mundhada1, Kevin S. Chou1, S. Harvey Moseley, Jr.1, Robert J. Schoelkopf1,2,3
${}^{\textrm{\scriptsize 1}}$ Quantum Circuits, Inc., New Haven, CT, USA
${}^{\textrm{\scriptsize 2}}$ Departments of Applied Physics and Physics, Yale University, New Haven, CT, USA
${}^{\textrm{\scriptsize 3}}$ Yale Quantum Institute, Yale University, New Haven, Connecticut, USA
${}^{\textrm{\scriptsize 4}}$ Current address: Microsoft Quantum
${}^{\textrm{\scriptsize 5}}$ Current address: Department of Condensed Matter Physics, Weizmann Institute of Science, Rehovot, Israel
$**{}^{\textrm{\scriptsize*}}$ Equal Contribution
VI Author contributions
@@
J.D.T. and S.O.M. contributed to the proposal and concept of the SWS gate. S.O.M. led the design of the dual-rail hardware architecture. J.O.Y led hardware deployment, experimental setup, and system validation. B.C. led the device fabrication. C.U.L. led the cavity design, preparation, and testing. K.S.C. led two-qubit gate testing, measurements, and analysis. N.M., T.N., P.L., B.H.E., A.Agrawal, A.N., K.S.C, J.O.Y., and J.D.T. contributed to data acquisition and data analysis. T.N., N.M., P.L., A. Agrawal, K.S.C., and J.D.T. contributed to the SWS gate calibration procedure. J.D.T. and K.S.C. contributed to numerical simulations of the SWS gate. P.L., N.M., J.O.Y., T.N., A. Agrawal, B.H.E., G.L., K.S.C., and N.K. contributed to system bring-up and calibration. S.O.M., R.C., S.M.F., T.Z.C., R.D.L., H.M., D.K.W., and J.O.Y contributed to quantum hardware design and simulation. E.L. B.B., M.C., C.F., M.R. contributed to mechanical design. S.O.M. and R.C. contributed to coupler design and flux-pump delivery. C.U.L., C.K., J.O.Y., B.H.E., and N.K. contributed to cavity preparation and testing. J.S., A.B., F.D., H.H., D.V.S. contributed to device assembly and integration. D.D., J.O.Y., M.Maxwell, and J.S. contributed to laboratory infrastructure. S.S, M.Mirzaei, D.S., B.G., B.C., H.L., and L.Y. performed device fabrication. M.Maxwell, E.R., A.C., S.R.D., R.J., R.P., and O.R designed, tested, and built the RF control system. J.O.Y., S.O.M., G. L., N.K. and R.A.S. contributed to RF wiring and implementation of the readout chain. J.C.C., P.L., N.M., D.K.W, K.S.C., and T.K. contributed to data analysis of randomized benchmarking results and supporting simulations. J.D.T., S.O.M., A.N, N.M., T.N., K.S.C., H.M., T.Z.C., N.K., and T.S. contributed to an earlier version of the SWS gate. M.S., S.W., D.Sweeney, A.Anderson, W.B., J.C., L.C., K.K., T.K., A.M., L.M-K., D.Szperka, I-W.S., Z.Y. contributed to software infrastructure and calibration tools used in this project. K.S.C, N.M., T.N., J.D.T., and R.J.S. wrote the manuscript. R.J.S., S.H.M., J.A., S.H., O.B. C-L.T. managed the project.
VII Materials and Correspondence
Correspondence and requests for materials should be addressed to Nitish Mehta (mehta@quantumcircuits.com), Kevin Chou (chou@quantumcircuits.com), and Robert J. Schoelkopf (robert.schoelkopf@yale.edu)
VIII Competing interests statement
@@
Supplemental information for “Bias-preserving and error-detectable entangling operations in a superconducting dual-rail system”
Appendix A Single-qubit characterization
A.1 1Q gate characterization